OpenStudy (anonymous):

Solve for x -32-3/5x=17

OpenStudy (anonymous):

is that -3/5x?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

-32-3/5x=17

OpenStudy (anonymous):

what can you do to get -3/5x on the left side by itself?

OpenStudy (anonymous):

+32

OpenStudy (anonymous):

Now I have -3/5x=15

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

imconfused to where you got 15? isnt your equation -32(-3/5x)=17 ?? +32 +32 (-3/5x) = 49 ? Did i read your equation wrong? haha sorry

OpenStudy (anonymous):

-17 sorry

OpenStudy (anonymous):

oohhhhhhhh hhahaha yes then in that case you are correct ! now that we are at (-3/5x) = 15